Katholieke Universiteit Leuven (KUL) Department of Computing Science

Heverlee Belgium

Researchers Hilde Adé, Maurice Bruynooghe, Luc Dehaspe, Luc De Raedt, Gunther Sablon, Wim Van Laer.

ILP Areas Theory revision, Multiple predicate learning, Autonomous agents, Knowledge discovery in databases.

Description of Research The research of the Leuven’s Machine Learning Group is concentrated on Theory Revision and clausal discovery in deductive databases. In the past few years, the interactive theory revisor CLINT [2] has been developed. This system inductively and abductively learns simple logic programs from examples and background knowledge. CLINT is a user-friendly and opportunistic multi-strategy learner that integrates several interesting features: it generates its own examples, shifts its language bias, learns multiple related concepts at the same time, employs integrity constraints, has a sound theoretical behaviour (it identifies concepts and knowledge bases in the limit), invents new concepts by analogy to previously learned ones and is able to reason about the unknown and the inconsistent using a special multi-valued logic. CLINT was the first concept-learner to use integrity constraints, to shifts the bias of first order logic concept description languages and is certainly one of the most advanced integrated inductive learning systems. Recently the development of non-interactive theory revision system RUTH [1] was started. This system updates a given theory (logic program), as to make it consistent with a given set of integrity constraints. For this purpose, several techniques that were used in CLINT, are adapted for a non-interactive setting.

CLAUDIEN [6] is the first ILP system that discovers full clausal theories from databases. It also uses a different semantics for ILP, which puts several issues in ILP into a new perspective. The development of CLAUDIEN is a result of cross-fertilization between the work on computational logic (and deductive databases), and ILP, which Leuven is involved in.

Recent research interests of Leuven include also the integration of inductive machine learning principles with planning. Current research involves the development of a planning system for an autonomous agent, that uses the logical representation of the event calculus [11, 12].

Recent Publications

[1] H. Adé, L. De Raedt, M. Bruynooghe. Theory revision. Proceedings of the Third International Workshop on Inductive Logic Programming, 1993.

[2] L. De Raedt. Interactive Theory Revision: An Inductive Logic Programming Approach. Academic Press, 1992.

[3] L. De Raedt, M. Bruynooghe. Interactive concept-learning and constructive induction by analogy. Machine Learning, Vol. 8, 107–150, 1992.

[4] L. De Raedt, M. Bruynooghe. Belief updating from integrity constraints and queries. Artificial Intelligence 53, 291–307, 1992.

[5] L. De Raedt, N. Lavrač. The many faces of inductive logic programming (invited talk). In J. Komorowski, Z. Ras (eds.) Proceedings of the 7th International Symposium on Methodologies for Intelligent Systems, Lecture Notes in Artificial Intelligence, Vol. 689, Springer-Verlag, 1993.

[6] L. De Raedt, M. Bruynooghe. A theory of clausal discovery. Proceedings of the 13th International Joint Conference on Artificial Intelligence, 1993.

[7] L. De Raedt. N. Lavrač, S. Džeroski. Multiple predicate learning. Proceedings of the 13th International Joint Conference on Artificial Intelligence, 1993.

[8] L. De Raedt. Inductive logic programming and scientific discovery, 1994 (in press).

[9] L. De Raedt. A brief introduction to inductive logic programming (tutorial). Proceedings of the International Logic Programming Symposium, The MIT Press, Vancouver, Canada, 1993.

[10] S. Muggleton, L. De Raedt. Inductive logic programming: Theory and methods. Journal of Logic Programming, 1994 (in press).

[11] G. Sablon, M. Bruynooghe. Using the event calculus to integrate planning and learning in an intelligent autonomous agent. Proceedings of the ECML ’93 Workshop on Integrated Learning Architectures, Vienna, 1993.

[12] G. Sablon, M. Bruynooghe. Using the event calculus to integrate planning and learning in an intelligent autonomous agent. Proceedings of the 2nd European Worskhop on Planning, Vadstena, Sweden, IOS Press, 1993.

Contact Persons Luc De Raedt Maurice Bruynooghe Department of Computing Science Celestijnenlaan 200 A B - 3001 Heverlee Belgium Phone: +32 16 20 10 15 Fax: +32 16 20 53 08 E-mail: {lucdr,maurice}@cs.kuleuven.ac.be